OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.

OpenStudy (mertsj):

What is the slope?

sam (.sam.):

\[slope=\frac{y-y_1}{x-x_1}\]

OpenStudy (seattle12345):

Isn't standard form A+B=C. While y= mx+b slope intercept form?

OpenStudy (anonymous):

oh yeah, your right!!

OpenStudy (mertsj):

Well just find the equation. You can put it in the correct form after you know what it is.

OpenStudy (mertsj):

What is the slope?

OpenStudy (anonymous):

Y=MX+B

OpenStudy (mertsj):

slope = change in y over change in x

OpenStudy (mertsj):

Subtract the y values of the ordered pairs. Subtract the x values of the ordered pairs. Make a fraction out of your answers.

OpenStudy (anonymous):

4 - 3 = 1 (y coordinate), 4 - 1 = 3 (x coordinate) y = 1/3x + b <---- slope intercept form (plug one of the coordinates in to solve for y intercept): 3 = 1/3(1) + b -1/3 -1/3 (algebraically manipulate) b (y intercept) = 8/3 or 2 and 2/3 y = 1/3x + 8/3 or -1/3x + y = 8/3 ---> Standard form Ax + B = C (Is that right, Mertsj?)

OpenStudy (mertsj):

4-3=1 4-1=3 slope = 1/3

OpenStudy (mertsj):

Yes Josh, that is right.

OpenStudy (anonymous):

Great :) thanks a lot, I thought i'd forgotten

OpenStudy (mertsj):

Good job.